OpenStudy (anonymous):

given p = 2 and t = 3, evaluate.p2-t2/4p+4t

OpenStudy (anonymous):

be aware that the numbers are exponents

OpenStudy (lgbasallote):

(2)^2 - (3)^2/ 4(2) + 4(3)

OpenStudy (lgbasallote):

what's the problem there?

jimthompson5910 (jim_thompson5910):

(p^2-t^2)/(4p+4t) ((2)^2-(3)^2)/(4(2)+4(3)) (4-9)/(4(2)+4(3)) (4-9)/(8+12) -5/20 -1/4 So (p^2-t^2)/(4p+4t) = -1/4 when p = 2 and t = 3

OpenStudy (anonymous):

how exactly did u get that answer? because i dont understand

jimthompson5910 (jim_thompson5910):

I plugged in p = 2 and t = 3. From there, I followed the order of operations to evaluate/simplify.